# TOP-Nav: Legged Navigation Integrating Terrain, Obstacle and Proprioception Estimation

### v2

@@ Abstract
Legged navigation is typically examined within open-world, off-road, and challenging environments. In these scenarios, estimating external disturbances requires a complex synthesis of multi-modal information. This underlines a major limitation in existing works that primarily focus on avoiding obstacles. In this work, we propose TOP-Nav, a novel legged navigation framework that integrates a comprehensive path planner with Terrain awareness, Obstacle avoidance and close-loop Proprioception. TOP-Nav underscores the synergies between vision and proprioception in both path and motion planning. Within the path planner, we present and integrate a terrain estimator that enables the robot to select waypoints on terrains with higher traversability while effectively avoiding obstacles. In the motion planning level, we not only implement a locomotion controller to track the navigation commands, but also construct a proprioception advisor to provide motion evaluations for the path planner. Based on the close-loop motion feedback, we make online corrections for the vision-based terrain and obstacle estimations. Consequently, TOP-Nav achieves open-world navigation that the robot can handle terrains or disturbances beyond the distribution of prior knowledge and overcomes constraints imposed by visual conditions. Building upon extensive experiments conducted in both simulation and real-world environments, TOP-Nav demonstrates superior performance in open-world navigation compared to existing methods. Project page at top-nav-legged.github.io.
I INTRODUCTION
Figure 1: Example deployment scenarios for TOP-Nav in both simulation and the real world. Besides obstacle avoidance, the robot plans an optimized direction on terrains with better traversability. For novel terrains, the robot incorporates proprioception history from previously traversed terrain to infer the traversability.
Imagine an outdoor hiking scenario where gait stability and efficient navigation are both critical. The target environment is often marked by intricate obstacles and hazardous terrains that cannot be fully observed through onboard vision. This challenge underscores the need for a comprehensive path planner integrating multi-modal observations. To achieve this, we humans employ experienced guidance and alternative perception modalities such as trekking poles and GPS devices. This example reveals a requirement of numerous trials and a substantial foundation of prior knowledge to accomplish the challenge navigation task.
@@ I INTRODUCTION
Although recent advancements in legged locomotion have allowed the robots to navigate various terrains based on a simulation-learned strong controller [21, 20, 5, 42, 40], the complexity of currently insimulable real-world factors makes it impossible for the robot to traverse all potential terrains encountered in reality. As a result, integrating the locomotion controller with only an open-loop path planner often restricts legged navigation to limited scenarios [18, 4, 23, 34].
@@
An effective solution to overcome these limitations is equipping the robot with terrain awareness. A traversable path can be planned based on the robot’s preferences on terrains [9, 15] with an appropriate distribution of contact heights and forces [8]. Unlike obstacle estimation, the distinct terrain features are typically encoded as semantic information, where traditional methods collect sufficient data and train segmentation or classification models [16] to learn these features. Nevertheless, compiling an exhaustive catalogue of all conceivable terrains along with their corresponding walking preferences is impractical [11]. Compounding the issue, the dynamic real-world conditions, such as lighting, humidity, and temperature, may introduce inaccuracies in the correspondence between images and walking preferences, especially when relying exclusively on vision in this context [39].
@@
The mentioned challenges can be attributed to the reliance on vision and the ignorance of motion states in path planning. To address this, we complement the vision-only terrain estimator with online corrections from motion evaluations. We construct a proprioception advisor to not only convey information about the traversability cost of novel terrains but also alert the robot to unexpected disturbances, such as invisible obstacles.
By integrating the Terrain estimator, Obstacle estimator, and Proprioception advisor, we formulate TOP-Nav: a hierarchical path and motion planning framework designed to navigate a quadruped robot through diverse and challenging terrains proficiently. TOP-Nav maintains four real-time robot-centric costmaps corresponding to: 1) goal approaching; 2) terrain traversibility; 3) obstacle occupancy; and 4) proprioception advice. The costmaps are synthesized with dynamically weighted factors to ensure a balanced consideration of safety and efficiency. We evaluate TOP-Nav both in simulation and on a physical robot, with a comparative analysis against existing approaches that address subsets of the factors (terrain, obstacle, proprioception). In simulation, we construct an environment featuring diverse terrains, including slopes, steps, and random textures, along with various obstacles. For real-world experiments, we deploy our approach in diverse off-road navigation scenarios encompassing common field terrains. Our main contributions are summarized below:
@@
II RELATED WORK
The pivotal feature of TOP-Nav lies in its integration of the terrain estimator and proprioception advisor. To elucidate this aspect, we present an overview of the relevant literature.
II-A Vision and Legged Proprioception Integration
Recent advancements in legged locomotion have showcased a synergistic mechanism for processing vision and proprioception within the context of “How to Walk” [2]. While heightmaps serve as crucial observations for a controller to generate dynamic motions across various terrains [27, 26], proprioception observations can be employed to reconstruct the heightmaps in visually degraded environments [30]. Different from the mentioned efforts focus on the depth channel, we propose a data-efficient solution to extract semantic information from the proprioceptive feedback.
The insights of integrating vision and legged proprioception are further explored in guiding the robot in “Where to Walk”. From the perspective of vision adied navigation, researchers have explored both hierarchical [23, 4] and end-to-end [32] pipelines. These methods require substantial efforts to develop a robust perception module [19], making it difficult and costly to transfer them to different hardwares.
@@ II-A Vision and Legged Proprioception Integration
To provide a comprehensive task observation and reduce the reliance on vision systems, recent researches have introduced proprioception to improve task planning. A majority of these works learn proprioception representations along with visual features in simulation, and then implement the cross-modal features through end-to-end [38], hybrid [17] or decoupled methods [41]. Despite the effectiveness demonstrated in these works, the high-dimensional representation space presents challenges for adaptation to novel scenarios and sim-to-real transfer. Alternatively, Fu et al. [14] introduced a hierarchical navigation framework that derives evaluation scores from motion states, yet it overlooks the integration of visual observations.
To mitigate these limitations, we propose a novel approach within TOP-Nav by maintaining a series of lightweight cost maps derived from multi-modal observations. This integration achieves a dynamic balance between vision and proprioception. Furthermore, we leverage the learning-based locomotion controller to derive motion evaluations from the value function, offering an efficient solution without additional training.
II-B Terrain Traversability Estimation
Terrain traversability is determined by factors such as terrain geometry, texture, and physical properties [12]. Estimating these features could be achieved by identifying the semantic class with a predefined static traversability score[31, 16, 35, 9]. These solutions exhibit a notable dependency on large-scale datasets [29] or limited to structured environments like urban scenarios [6, 1].
@@ II-B Terrain Traversability Estimation
In off-road navigation, the motion states involved in the dynamic interactions between the robot and the environment provide valuable metrics for assessing terrain traversability [10]. These insights have inspired methods that eliminate the need for manual annotation by autonomously deriving terrain traversability from proprioception through self-supervised learning [39, 3, 22, 24, 7, 28]. Nevertheless, the performance of these studies is contingent upon the quality of the collected datasets [12]. To emphasize the challenges in unconstrained navigation, researchers have proposed various approaches to handling novel observations. For instance, Frey et al. [11] updated the traversability estimation network online with anomalies into consideration. Karnan et al. [25] performs nearest-neighbor search in the proprioception space to align visually novel terrains with existing traversability.
Drawing inspiration from those works estimating traversability for novel terrains, we propose a prior-knowledge informed terrain estimator that employs the proprioception advisor as online corrections. Our method diverges from previous approaches primarily in two key aspects: 1) We employ an estimated value function from reinforcement learning to assess terrain traversability, providing a comprehensive evaluation of robot-terrain interactions. 2) By incorporating the proprioception based terrain traversability estimation as online corrections to vision-based predictions, we achieve a more data-efficient approach for estimating traversability on unknown terrains compared to online training.
III Background
The proposed proprioception advisor leverages recent progress in legged locomotion to provide online motion evaluations for the path planner. This section provides an essential background of this.
@@ III Background
Learning-based legged locomotion controllers has been well developed through reinforcement learning, which is generally achieved by updating the policy ${\pi}$ within the asymmetric actor-critic training:
@@
$L^{\textrm{critic}}_{t}=(c_{t}-c_{t}^{\textrm{targ}})^{2},$ (2)
where
$c_{t}^{\textrm{targ}}=\sum_{i=t}^{T}\gamma^{i-t}R(s_{i}),$ (3)
@@
${s_{i}}$ denotes the robot state, $R(s_{i})$ represents the rewards accrued at timestep $i$, which commonly includes guiding the robot to track a given velocity commands with stable gaits and attitude. A substantial efforts in reward engineering has enabled the value function $c_{t}^{\textrm{targ}}$ to evaluate a comprehensive set of interactions between the robot and its environment. Therefore, when deploying the learning-based controller, $c_{t}$ can be utilized to provide closed-loop feedback for the task planner.
@@
The complete training paradigm will involve a second stage, training a depth encoder and a student network to reproducing ${\bm{o}_{t}^{p}}$ and ${\bm{o}_{t}^{e}}$ from real world accessible observations $\bm{I}_{d}$ (depth image), $\bm{o}_{t}^{i}$ and $\bm{o}_{t}^{h}$. The motion controller will track the planned $\Delta_{\textrm{yaw}}$ and $v_{\textrm{lin}}$ and the control action is represented by the 12 desired joint position $\bm{q}_{\textrm{des}}$.
Figure 2: TOP-Nav: the hierarchical path and motion planning framework to tackle the problem of legged navigation with a point goal. The path planner synthesizes relative distance, terrain traversability, obstacle occupancy, and motion evaluations into a combined cost map, from which it computes waypoints based on the overall cost considerations. The motion evaluations are extracted from a learning-based motion planner, which receives depth observation as input.
IV Method
IV-A System Overview
As illustrated in Fig. 2, the proposed TOP-Nav pipeline connects a path and motion planner to tackle the task of legged navigation. To efficiently generate a collision-free path and stable motion states, we highlight a terrain estimator and a proprioception advisor inside the pipeline. In addition to combining these modules on the costmap (Fig. 2), the proprioception advisor provides online corrections for the terrain estimation. Such integration informs the robot of the traversability of unknown terrains and invisible obstacles, allowing TOP-Nav to handle open-world task without being constrained by visual conditions or prior knowledge.
To elucidate the task configuration, the robot is given a point goal with its location $\bm{p}_{\textrm{goal}}$ and has to approach the goal within a limited time. The location of the robot is denoted as $(\bm{p}_{\textrm{base}},r_{\textrm{base}})$. For external observation, we utilize a bio-channel perception module for both the task and motion planner. This module includes depth ($\bm{I}_{d}$) and RGB ($\bm{I}_{\textrm{rgb}}$) channels. The path planner computes the desired velocity command $(v_{\textrm{lin}},\Delta_{\textrm{yaw}})$ for the robot, which is tracked by the motion planner along with the appropraite joint position signal.
IV-B Integrated Path Planner

### v3

@@ Abstract
Legged navigation is typically examined within open-world, off-road, and challenging environments. In these scenarios, estimating external disturbances requires a complex synthesis of multi-modal information. This underlines a major limitation in existing works that primarily focus on avoiding obstacles. In this work, we propose TOP-Nav, a novel legged navigation framework that integrates a comprehensive path planner with Terrain awareness, Obstacle avoidance and close-loop Proprioception. TOP-Nav underscores the synergies between vision and proprioception in both path and motion planning. Within the path planner, we present a terrain estimator that enables the robot to select waypoints on terrains with higher traversability while effectively avoiding obstacles. In the motion planning level, we construct a proprioception advisor from the learning-based locomotion controller to provide motion evaluations for the path planner. Based on the close-loop motion feedback, we offer online corrections for the vision-based terrain and obstacle estimations. Consequently, TOP-Nav achieves open-world navigation that the robot can handle terrains or disturbances beyond the distribution of prior knowledge and overcomes constraints imposed by visual conditions. Building upon extensive experiments conducted in both simulation and real-world environments, TOP-Nav demonstrates superior performance in open-world navigation compared to existing methods. Project page at top-nav-legged.github.io.
Keywords: Navigation, Task Planning, Reinforcement Learning
1 Introduction
Figure 1: TOP-Nav achieves open-world navigation in both simulation and the real world. The robot plans an obstacle-free path on terrains with better traversability. For novel terrains, the robot rapidly estimates its traversability based on proprioception experience.
@@ 1 Introduction
Despite recent advancements in legged locomotion allowing the robots to navigate various terrains based on a simulation-learned strong controller [1, 2, 3, 4, 5, 6, 7, 8, 9], the complexity of currently insimulable real-world factors makes it impossible for the robot to traverse all potential terrains encountered in reality. As a result, integrating the locomotion controller with only an open-loop path planner often restricts legged navigation to limited scenarios [10, 11, 12, 13].
@@
An effective solution to overcome these limitations is equipping the robot with terrain awareness: a traversable path can be planned based on the robot’s preferences on terrains [14, 15] with an appropriate distribution of contact heights and forces [16]. Unlike obstacle estimation, the distinct terrain features are typically encoded as semantic information, where traditional methods collect sufficient data and train segmentation or classification models [17] to learn these features. Nevertheless, compiling an exhaustive catalogue of all conceivable terrains along with their corresponding walking preferences is impractical [18]. Compounding the issue, the dynamic real-world conditions, such as lighting, humidity, and temperature, may introduce inaccuracies in the correspondence between images and walking preferences, especially when relying exclusively on vision in this context [19].
@@
To address the mentioned challenges that relying solely on vision and ignoring motion states in path planning, we complement the vision-only terrain estimator with online corrections derived from motion evaluations. We construct a proprioception advisor to not only convey information about the traversability cost of novel terrains but also alert the robot to unexpected disturbances, such as invisible obstacles.
By integrating the Terrain estimator, Obstacle estimator, and Proprioception advisor, we formulate TOP-Nav: a hierarchical path and motion planning framework navigating a quadruped robot through diverse and challenging terrains proficiently. We evaluate TOP-Nav both in simulation and on a physical robot, with a comparative analysis against existing legged navigation systems. Our main contributions are summarized below:
@@
2 RELATED WORK
2.1 Vision and Legged Proprioception Integration
@@
Vision-aided legged navigation has been extensively explored in existing literature [12, 11, 20], with performance heavily dependent on a robust perception module [21]. This reliance makes it difficult and costly to transfer a specific system to different hardware platforms. To provide a comprehensive task observation and reduce the reliance on vision systems, recent researches have introduced proprioception to improve task planning. A majority of these works learn proprioception representations along with visual features in simulation, and then implement the cross-modal features through end-to-end [22], hybrid [23] or decoupled frameworks [24]. Despite the effectiveness demonstrated in these works, the high-dimensional representation space presents challenges for adaptation to novel scenarios and sim-to-real transfer. Alternatively, Fu et al. [25] introduced a hierarchical navigation framework that derives evaluation scores from motion states, yet it overlooks the integration of visual observations. To mitigate these limitations, we propose a novel approach within TOP-Nav by maintaining a series of lightweight cost maps derived from multi-modal observations. This integration achieves a dynamic balance between vision and proprioception. Furthermore, we leverage the learning-based locomotion controller to derive motion evaluations from the value function, offering an efficient solution without additional training.
2.2 Terrain Traversability Estimation
@@
Terrain traversability is determined by factors such as terrain geometry, texture, and physical properties [26]. Estimating these features could be achieved by identifying the semantic class with a predefined static traversability score[27, 17, 28, 14]. These solutions exhibit a notable dependency on large-scale datasets [29] or limited to structured environments like urban scenarios [30, 31]. In off-road navigation, the motion states involved in the dynamic interactions between the robot and the environment provide valuable metrics for assessing terrain traversability [32]. These insights have inspired methods that eliminate the need for manual annotation by autonomously deriving terrain traversability from proprioception through self-supervised learning [19, 33, 34, 35, 36, 37]. Nevertheless, the performance of these studies is contingent upon the quality of the collected datasets [26]. To emphasize the challenges in unconstrained navigation, researchers have proposed various approaches to handling novel observations. For instance, Frey et al. [18] updated the traversability estimation network online with anomalies into consideration. Karnan et al. [38] performs nearest-neighbor search in the proprioception space to align visually novel terrains with existing traversability. Drawing inspiration from those works estimating traversability for novel terrains, we propose a terrain estimator that employs the proprioception advisor as online corrections. Our method diverges from previous approaches primarily in two key aspects: 1) We employ an estimated value function from reinforcement learning to assess terrain traversability, providing a comprehensive evaluation of robot-terrain interactions. 2) We make online corrections on the vision-based estimation without additional training, providing a data-efficient solution for identifying novel terrains.
3 Background
@@
Learning-based legged locomotion controllers have been well developed through reinforcement learning, which is generally achieved by updating the policy ${\pi_{1}}$ within the asymmetric actor-critic training:
@@
$L^{\textrm{critic}}_{t}=(c_{t}-c_{t}^{\textrm{targ}})^{2}\text{ and }c_{t}^{\textrm{targ}}=\sum_{i=t}^{T}\gamma^{i-t}R(s_{i}),$ (2)
@@
${s_{i}}$ denotes the robot state, $R(s_{i})$ represents the rewards accrued at timestep $i$, which commonly includes guiding the robot to track a given velocity commands with stable gaits and attitude. Substantial efforts in reward engineering to formulate $R(s_{i})$ have enabled the value function $c_{t}^{\textrm{targ}}$ to evaluate a comprehensive set of interactions between the robot and its environment. This provides an essential foundation for the proposed proprioception advisor to leverage the estimated $c_{t}$ for motion evaluations within the path planner.
@@
The complete training paradigm will involve a second stage, training a depth encoder and a student network to reproducing ${\bm{o}_{t}^{p}}$ and ${\bm{o}_{t}^{e}}$ from real world accessible observations $\bm{I}_{d}$ (depth image), $\bm{o}_{t}^{i}$ and $\bm{o}_{t}^{h}$. The motion controller will track the target direction $\Delta_{\textrm{yaw}}$ and velocity command $v_{\textrm{lin}}$, the control signal is represented by the desired joint position $\bm{q}_{\textrm{des}}$.
4 Method
4.1 System Overview